# Reduced Order Modelling, Dynamical Systems, Deep Learning, Uncertainty Quantification journal: Journal name

### v2

@@ 6 Conclusion
This paper presents a robust and parametrised reduced-order model (ROM) for time-dependent partial differential equations applied to the Navier-Stokes equations in the laminar regime. The model makes use of memory embedding and variational inference, first using a nonlinear variational autoencoder to reduce the high-dimensional dynamics to a low-dimensional latent space, and second using an attention-based transformer to capture time dependencies in the reduced manifold. In addition, cross-attention incorporates dependencies on external parameters. This allows UP-dROM to handle different dynamical regimes, making it adaptable to external excitation variables and enhancing generalisation.
@@
Our model demonstrates high accuracy, with a relative mean square error consistently below 0.5% on the test set for seen parameters. Interpolation between seen parameters remains below 2.5%, even when trained with minimal parameter variety. Extrapolation is effective in some regions, but less reliable in others. However, the model’s Uncertainty Quantification (UQ) framework allows these uncertain regions to be identified a priori. A key contribution of this work is the integration of a computationally efficient UQ framework that provides instantaneous confidence estimates without increasing the computational burden during inference. The approach links uncertainty to both system variability and distance from the training distribution, providing an interpretable measure of confidence across time, spatial domains or parameter space.
@@
We show that UQ is not only valuable in itself, but also a useful tool for efficiently sampling the parameter space. As it is often impractical to collect training data for the entire domain, training is restricted to a small subset of parameters. Furthermore, we show a strong positive linear correlation between the error and the uncertainty in the parameter space. Therefore, quantifying the uncertainty of the model can guide adaptive sampling, focusing retraining efforts on regions of high uncertainty. This reduces the need for validation data across the entire domain, optimising learning and enhancing generalisation across different dynamical regimes. Furthermore, the framework effectively captures bifurcation dynamics and complex parametric dependencies, making it well suited for modelling nonlinear fluid systems.
@@
Looking forward, future work will explore data assimilation strategies to refine the model in uncertain regions. The model can identify regions of high uncertainty in physical space where sparse sensor data can be used to improve the model without the need for additional full-state training data. Another promising direction is the application of the model to fluid dynamics interacting with moving structures. In addition, using the model’s parametrisation as a control variable within a control framework could open up new possibilities for flow optimisation applications.

### v3

@@ 6 Conclusions and perspectives
This paper presents a robust and parametrised reduced-order model (ROM) for time-dependent partial differential equations applied to the Navier-Stokes equations in the laminar regime. The model uses memory embedding and variational inference, first, a nonlinear variational autoencoder is used to reduce the high-dimensional dynamics to a low-dimensional latent space and second, an attention-based transformer captures time dependencies in the reduced manifold. In addition, cross-attention incorporates dependencies on external parameters. This allows UP-dROM to handle different dynamical regimes, making it adaptable to external excitation variables and enhancing generalisation.
@@
Our model demonstrates high accuracy, with a relative mean square error consistently below 0.5% on the test set for seen parameters. Interpolation between seen parameters remains below 2.5%, even with limited parameter diversity. Extrapolation is effective in some regions, but less reliable in others. However, the model’s Uncertainty Quantification (UQ) framework allows these uncertain regions to be identified a priori. A key contribution of this work is the integration of a computationally efficient UQ framework that provides instantaneous confidence estimates without adding to the computational cost during inference. The approach links uncertainty to both system variability and distance from the training distribution, providing an interpretable measure of confidence across time, spatial domain, and parameter space.
@@
We show that UQ is not only valuable in itself, but also a useful tool for efficiently sampling the parameter space. Since collecting training data across the entire domain is often impractical, training is restricted to a small subset of parameters. Furthermore, we show a strong positive linear correlation between the error and the uncertainty in the parameter space. Therefore, quantifying the uncertainty of the model can guide adaptive sampling, focusing retraining efforts on regions of high uncertainty. This reduces the need for validation data across the entire domain, optimising learning and enhancing generalisation across different dynamical regimes. In addition, the framework effectively captures bifurcation dynamics and complex parametric dependencies, making it well suited for modelling nonlinear fluid systems.
@@
Future work will explore data assimilation strategies to refine the model in uncertain regions. The model can identify regions of high uncertainty in physical space where sparse sensor measurements can enhance the model without requiring additional full-state training data. In addition, using the model’s parametrisation as a control variable within a control framework could open new possibilities for flow optimisation.